ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

3. Number Theory and Arithmetic Algebraic Geometry

1991 MS Classification: 11D41 Tirado, Eleazar, Apartado 80890, Caracas 1080 A, Venezuela Three arguments that conform a basis for an approach to Fermat’s last theorem On how the concept of proper Fermat’s equations, Shank’s theorem 48 and the parametric equations through which two terms of the equation, (x, y) → (x^{p}, y^{p}) with p (prime)≥ 3, may attain all those values such that (x ± y) = N^{p} : N ∈ ZZ conform a basis onto which an approach to Fermat’s last theorem can be formulated.